OpenStudy (anonymous):

anyone good at synthetics? from alg 2 Help anyone?

OpenStudy (anonymous):

please help me

jimthompson5910 (jim_thompson5910):

do you have a specific question?

OpenStudy (anonymous):

ya its (2t^2+13t+15) divided by (t+5)

jimthompson5910 (jim_thompson5910):

Start by writing out the coefficients in a table and write the test root (-5) off to the left like this -5 | 2 13 15 | ------------------------------

jimthompson5910 (jim_thompson5910):

Pull down the first coefficient 2 -5 | 2 13 15 | ------------------------------ 2

jimthompson5910 (jim_thompson5910):

Multiply the test root (-5) with 2 to get -10, write this under the 13 -5 | 2 13 15 | -10 ------------------------------ 2

OpenStudy (anonymous):

i did that

jimthompson5910 (jim_thompson5910):

Add 13 to -10 to get 3. Write this under the -10 -5 | 2 13 15 | -10 ------------------------------ 2 3

jimthompson5910 (jim_thompson5910):

Repeat the last two steps -5 | 2 13 15 | -10 -15 ------------------------------ 2 3 0

jimthompson5910 (jim_thompson5910):

The last line is 2 3 0 so (2t^2+13t+15) divided by (t+5) = 2t+3 remainder 0

jimthompson5910 (jim_thompson5910):

Since the remainder is 0, t+5 is a factor of 2t^2+13t+15 ie 2t^2+13t+15 = (t+5)(2t+3)

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

can you help me with a few more problems if i get stuck?

jimthompson5910 (jim_thompson5910):

sure just a few though

OpenStudy (anonymous):

ok thanks